OpenStudy (anonymous):

Find the coordinates of the midpoint of the segment with the given points. J (-7,-5) and K (-3,7).

OpenStudy (anonymous):

take the average in each coordinate, i.e. add and divide by 2

OpenStudy (anonymous):

\[\left(\frac{-7-3}{2},\frac{-5+7}{2}\right)\] is a start